OpenStudy (anonymous):

sqrtx(3x+0), what is the derivative?

OpenStudy (anonymous):

please show calc

OpenStudy (anonymous):

\[\sqrt{x}(3x+0)\]

OpenStudy (anonymous):

ive got 9/2x ^1/2

OpenStudy (anonymous):

You want the derivative of that or the other one?

OpenStudy (anonymous):

i want the derivative of the entire f(x)=\[\sqrt{x}(3x+0)\]

OpenStudy (anonymous):

\[3x^{3/2}\] Derivative: \[\frac{3}{2}3x^{1/2}=\frac{9}{2}\sqrt{x}\]

OpenStudy (anonymous):

can you explain how you did it pls

OpenStudy (anonymous):

product rule = uv'+vu' u=sqrtx or x^(1/2) u'=.5x^(-1/2) v=3x v'=3 (x^(1/2))(3)+(3x)(.5^(-1/2)) simplify

OpenStudy (anonymous):

I've got the same answer as you, you were right

OpenStudy (anonymous):

rickjbr, my teacher wrote that answer on the board. i just didnt understand how he arrived there

OpenStudy (anonymous):

I just distribute the sqrt(x) to the parentheses and took the derivative of that

OpenStudy (anonymous):

\[\sqrt{x}(3x+0)=3(x)^{3/2}\] Then take the derivative: Multiply the exponent by the constant and subtract 1 from the exponent: Derivative: \[\frac{3}{2}*3x^{1/2}\]

OpenStudy (anonymous):

\[\frac{3}{2}3x^{3/2-2/2}=\frac{9}{2}\sqrt{x}\]

OpenStudy (anonymous):

thanks